### v1

# MindGYM: Enhancing Vision-Language Models via Synthetic Self-Challenging Questions

Abstract
Large vision-language models (VLMs) face challenges in achieving robust, transferable reasoning abilities due to reliance on labor-intensive manual instruction datasets or computationally expensive self-supervised methods. To address these issues, we introduce MindGYM, a framework that enhances VLMs through synthetic self-challenging questions, consisting of three stages: (1) Seed Single-Hop Question Synthesis, generating cognitive questions across textual (e.g., logical deduction) and multimodal contexts (e.g., diagram-based queries) spanning eight semantic areas like ethical analysis; (2) Challenging Multi-Hop Question Synthesis, combining seed questions via diverse principles like bridging, visual-textual alignment, to create multi-step problems demanding deeper reasoning; and (3) Thinking-Induced Curriculum Fine-Tuning, a structured pipeline that progressively trains the model from scaffolded reasoning to standalone inference. By leveraging the model’s self-synthesis capability, MindGYM achieves high data efficiency (e.g., +16% gains on MathVision-Mini with only 400 samples), computational efficiency (reducing both training and inference costs), and robust generalization across tasks. Extensive evaluations on seven benchmarks demonstrate superior performance over strong baselines, with notable improvements (+15.77% win rates) in reasoning depth and breadth validated via GPT-based scoring. MindGYM underscores the viability of self-challenging for refining VLM capabilities while minimizing human intervention and resource demands. Code and data are released to advance multimodal reasoning research.
**footnotetext: Equal contributions.
1 Introduction
Large vision-language models (VLMs) have emerged as foundational tools for tasks requiring joint reasoning over visual and textual inputs [4, 30]. However, enabling robust performance with transferable and efficient reasoning abilities with high-quality data remains challenging [6, 8]. Manually curated instruction datasets like OK-VQA [25] and ScienceQA [22] are labor-intensive to scale, while self-supervised synthetics methods such as MMInstruct [21] and MMEvol [24] suffer from limited generalization cross modality and tasks, failing to tightly and logically couple visual and textual semantics. Meanwhile, reasoning enhancement methods—such as reinforcement learning (RL) [1] or iterative prompting [19]—incur prohibitive computational costs, limiting their practicality.
To address these challenges, we propose MindGYM, a new framework illustrated in Figure 1 that enhances VLMs via synthetic self-challenging questions with three core stages: (1). Seed Single-Hop Question Synthesis: Generates cognitive questions across text-only (e.g., logical deduction) and multimodal contexts (e.g., scientific diagrams), spanning eight semantic topics like causal reasoning and ethical analysis. (2). Challenging Multi-Hop Question Synthesis: Combines seed questions using cognitive operators—textual strategies (Bridging, Comparison, Temporal) and multimodal reasoning types (Visual-Textual Alignment, Spatial Reasoning, etc.)—to create layered, multi-step problems demanding deeper cognitive engagement. (3). Thinking-Induced Curriculum Fine-Tuning: A structured training pipeline progresses from scaffolded reasoning (e.g., answer prediction with explicit thought paths) to standalone inference, embedding advanced reasoning capacities into model parameters. All these processes are operated on the model to be tuned, via cognitive data self-synthesis and self-challenging training.
Extensive empirical evaluations across 7 benchmarks demonstrate that MindGYM consistently outperforms prior methods on both text-based and multimodal scenarios, achieving +16% gains on challenging tasks like MathVision-Mini with merely 400 synthesis samples, while reducing computational costs in both training and inference time. Notably, MindGYM gains robust generalization across different synthesis sources, and enhanced reasoning depth and breadth via GPT-based scoring analysis over state-of-the-art (SOTA) competitors (+15.77% win rates), enabling the model to handle open-ended tasks requiring multi-modal, high-order cognition.
Our main contributions can be summarized as follows:
  • We propose a new cognitive-grounded synthetic data synthesis pipeline, that systematically generates complex questions considering thinking depth, thinking width, and cross-modality.
  • We show the feasibility of self-challenging evolution for VLMs, with high data-efficiency and progressive training that absorbs explicit human-like reasoning patterns.
  • We verify MindGYM’s effectiveness, practical efficiency, and contextual adaptability with extensive experiments. Our code and data are released at https://github.com/modelscope/data-juicer/tree/MindGYM to foster future progress in vision-language reasoning.
Figure 1: The proposed MindGYM framework architecture consists of three core stages: ① Seed Single-Hop Question Synthesis, ② Challenging Multi-Hop Question Synthesis, and ③ Thinking-Induced Curriculum Fine-Tuning. The process begins with cognitive-aware self-synthesis across text (e.g., mathematical/biological questions) and multimodal (text+image) contexts. Seed questions are then combined via text-based strategies (Bridging, Comparison, Temporal) and multimodal reasoning types (Visual-Textual Alignment, Spatial Reasoning, Causal Inference, Contextual Synthesis) to form self-challenging questions. Finally, a curriculum-driven fine-tuning phase injects structured cognitive competence into the model through masked-task learning. Color-coded regions and arrows illustrate the hierarchical progression from simple reasoning to advanced problem-solving, emphasizing conceptual depth and modal integration.
2 Related Works
Instruction Data Construction for VLMs.
Large vision-language models (VLMs) critically depend on high-quality instruction-tuning datasets to achieve robust performance [imgdiff, 7]. Early efforts, such as OK-VQA [25] and ScienceQA [22], rely on human-annotated data to ensure accuracy and relevance. However, these manual approaches are resource-intensive and difficult to scale. To address this limitation, recent works have explored self-instruct paradigms that automate dataset creation. While text-centric methods like [39, 34, 42] emphasize diversity and complexity in textual instructions, their extension to multimodal domains remains challenging. Multimodal self-instruction frameworks, such as MMInstruct [21] and MMEvol [24], attempt to bridge this gap by iteratively generating image-text pairs. Nevertheless, these methods often suffer from modality inconsistency, where generated textual instructions lack tight visual grounding or rely on shallow cross-modal alignments (e.g., generic image descriptions). This issue arises due to an over-reliance on LLMs for text generation, which lack explicit mechanisms for visual reasoning. Consequently, the resulting datasets fail to challenge VLMs on fine-grained vision-language understanding tasks.
Reasoning-Aware Question Synthesis.
Synthetic question generation has shown success in domains like mathematical reasoning [20, 36], where structured problem spaces enable systematic scaling of complexity [32] and skill decomposition [3]. However, extending these techniques to multimodal domains poses unique challenges due to three key factors: (1) Heterogeneous Semantics: Vision-language reasoning requires handling unstructured inputs with diverse modalities, unlike the formal grammars available in mathematical problems. (2) Cross-Modal Synergy: Current methods focus on single-modality skill progression rather than fostering interactions between modalities. (3) Open-Ended Challenges: Template-driven approaches, such as Visual Program Induction [12], struggle to generate open-ended reasoning tasks that demand higher-order cognitive skills. As a result, existing methods exhibit limited generalizability to complex vision-language tasks, underscoring the need for more flexible and domain-agnostic synthesis frameworks.
Reasoning-Aware Self-Improvement.
Enhancing reasoning capabilities in VLMs often incurs significant computational overhead. RL-based approaches [1] require extensive policy optimization over large-scale reasoning trajectories [14], making them computationally expensive. Iterative refinement techniques [19] further exacerbate costs by introducing multi-stage processing pipelines. Even parameter-efficient methods, such as Chain-of-Thought prompting [40], suffer from latency penalties due to lengthy sequential decoding. Recent advances in test-time scaling [26] demonstrate promising performance gains but at the expense of increased inference latency. These limitations highlight the need for more efficient strategies to improve reasoning without compromising computational feasibility.
Our Position.
MindGYM addresses these three critical limitations of prior art: (1) Modality-inconsistency. With multiple cognitive relationship enhancement in multi-hop questions, we enforce tight semantic coupling in vision-language concepts. (2) Generalization. By adopting domain-agnostic prompts, we explicitly incorporate eight vision meta-topics beyond math, and adaptively increase visual grounding and linguistic complexity on the synthesis data. (3) Computational efficiency. By injecting complex, self-challenging questions directly into the training data, we encourage models through curriculum fine-tuning to perform deep and wide reasoning within a single inference step, eliminating the need for iterative adjustments.
3 The Proposed MindGYM
In this work, we aim to embed advanced reasoning capabilities into VLMs through synthesized fine-tuning datasets. For large models, conventional single-hop question-answer pairs are insufficient to stimulate reasoning in both depth and breadth. To address this, we propose a progressive self-synthesis framework, outlined in Figure 1, operating through three interconnected stages:
(1) Seed Question Synthesis: Starting with a base VLM $\pi$ (e.g., Qwen-2.5-VL [4]), the model self-generates logically related single-hop questions $\{q^{1}_{i},\ldots,q^{k}_{i}\}_{i\in N}$ to explore diverse reasoning methodologies and topics. Here, $k$ denotes the number of seed questions per context and $N$ represents the number of synthesized data.
(2) Self-Challenging Question Synthesis: Using $\{q^{1}_{i},\ldots,q^{k}_{i}\}$ as inputs, the model constructs structured complex questions $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$ via multi-hop combinations. Each $Q_{i}$ is formed as $Q_{i}=q^{1}_{i}\oplus\ldots\oplus q^{k}_{i}$, where $\oplus$ represents cognitive operators (e.g., composition, contradiction, abstraction) that create hierarchical reasoning structures. The model concurrently generates explicit thinking trajectories $T_{i}$, which documents the rationale behind $Q_{i}$ and $A_{i}$.
(3) Self-Challenging Cognitive Training: The model is fine-tuned on four structured datasets $\{D^{j}_{\text{train}}\}_{j=1}^{4}$ through a curriculum pipeline. Training focuses on predicting masked components across questions $(Q_{i})$, answers $(A_{i})$, and their associated thought paths $(T_{i})$, thereby cultivating deeper and broader reasoning capacity. This adversarial questioning approach systematically improves the model’s ability to navigate ambiguous, multi-step reasoning tasks.
The proposed framework enhances generalization by exposing models to diverse cognitive challenges while injecting interpretable thinking patterns that mirror human-like problem-solving dynamics into the model parameters. Detailed design and implementations are provided in subsequent subsections.
3.1 Seed Single-Hop Question Synthesis
3.1.1 Diverse Thinking Methodologies
Cognitive-Aware Synthesis Templates.
We adopt a synthesis template that injects thinking elements, consisting of three parts: cognitive setting, background conditions, and synthesis goals: $P=\{P_{cog},P_{context},P_{task}\}$. $P_{context}$ and $P_{task}$ have different instantiations in different synthesis scenarios (single-hop, multi-hop QA, text, multimodal) in our method  [^1] The full prompt is presented in the Appendix B., while $P_{cog}$ remains unchanged to encourage the model to consistently produce responses that are strongly related to thinking.
We first describe the single-hop question scenario, with other tasks detailed later. Based on the given $P_{context}$, $P_{task}$ requires $\pi$ to generate $k$ logically connected seed single-hop questions $q^{1}_{i},...q^{k}_{i}$, at the same time, the model will be explicitly noted that these seed questions will be used later as more complex multi-hop combinations (the specific combination types will be detailed in later Sec.3.2).
Adaptive Stream of Consciousness.
We draw inspiration from a thinking protocol named Thinking-Claude [44] and adapt it to shape $P_{cog}$ as a collection of many diverse thinking processes, (e.g., making and challenging hypotheses, exploring and balancing different perspectives). This makes the model mimic the human thinking process in a layered, context-aware style, aiming for a broader coverage in thinking methodologies. Given different $P_{context}$ and $P_{task}$, $\pi$ dynamically adjusts thinking dimensions, such as problem depth, interpretation angles, and validation rigor, and explicitly expose models’ “minds” dealing with complex scenarios. This adaptive exploration not only encourages the model to incorporate diverse reasoning paths for generalizing to reasoning tasks, but also injects structured variability into training data, countering the memorization biases in conventional pipelines. Empirical validation of $P_{cog}$ is presented in Section 4.3.
3.1.2 Diverse Thinking Objects
Meta-Topics.
We consider both text-only and text-image QA synthesis. Here we first introduce the $P_{context}$ for the text-only case. Building on the cognitive architecture theory [28] and the multi-compartmental knowledge representation [33], we establish eight orthogonal meta-topics spanning four foundational reasoning dimensions: quantitative, causal, temporal, and social-ethical. Unlike prior works focused on limited domain combinations [2], our approach emphasizes broad cognitive coverage:
  • Mathematical Reasoning: Solving quantitative problems using numbers and formulas.
  • Scientific Knowledge: Understanding natural laws and scientific principles.
  • Logical Deduction: Forming conclusions through logical progression from premises.
  • Technical Procedures: Following step-by-step instructions for practical tasks.
  • Historical Events: Analyzing past events and their consequences.
  • Ethical Considerations: Evaluating decisions through moral frameworks.
  • Economic Principles: Studying resource allocation and valuation.
  • Psychological Insights: Exploring human behavior and cognition.
These eight topics are selected for two core reasons: (1) alignment with widely studied reasoning domains (e.g., mathematics, logic) [17], and (2) extension to synergize with VLMs’ capabilities and potential [30], balancing foundational reasoning with subsequent self-challenges.
By polling one of these meta-topics and their descriptions, we require $\pi$ to self-generate background document as $P_{context}$, then self-generate $k=5$ seed single-hop questions per inference. Empirical results in Section 4.3 demonstrate that this topic-balanced sampling strategy enhances thinking breadth and significantly improves end-to-end model performance compared to unbalanced alternatives. This process allows $\pi$ to explore maximal topic diversity, leveraging its internal knowledge to identify domains and skills requiring advanced reasoning.
Anchoring Image Sources.
For the $P_{context}$ component in text-image QA synthesis, we use existing images rather than generating images from scratch (as in the text-only case). This is because the existing SOTA VLMs are mostly only image-to-text generation (i.e., image understanding) and fail to do text-to-image generation at the same time. This decision arises because current SOTA VLMs predominantly specialize in image-to-text generation (i.e., image understanding) and lack dual capabilities for text-to-image generation constrained by their architectures. Without anchored visual references, the model risks hallucinatory reasoning about image elements (e.g., referencing “growth curves” might lead to incorrect inferences about unrelated scenarios, such as spatial-based reasoning tasks).
Specifically, we randomly sample image-text pairs from scientific knowledge and general-domain QA datasets – ScienceQA [22] and OKVQA [25] – to serve as background context ($P_{\text{context}}$) for $\pi$. We emphasize alignment between visual and textual concepts within the generation instruction $P_{\text{task}}$, prompting $\pi$ to self-generate $k=5$ seed questions per inference. This enhances harmonized interactions between image and text modalities during synthesis.
Reject Sampling for Diversity.
To enhance diversity in seed question generation, MindGYM further applies semantic vectorization to textual components of data in the synthesis pool. We use an empirically determined cosine similarity threshold to evaluate the overlap between newly generated and existing samples. If similarity exceeds this threshold, the system initiates a regeneration cycle to eliminate semantically redundant content, continuing until the seed question pool reaches the target size $N$. This mechanism widens the breadth of reasoning in synthetic data, better preparing it for diverse downstream scenarios.
Unlike prior methods concentrating on domain-specific reasoning (e.g., mathematical tasks), MindGYM expands reasoning scope across three dimensions: cognitive-aware thinking methodologies, image-text interactions, and semantic redundancy filtering. In Section 4.2, we empirically validate MindGYM ’s robustness and effectiveness across varied synthesis sources—including model’s internal knowledge, multilingual contexts, and anchored visual content.
3.2 Challenging Multi-Hop QA Synthesis
3.2.1 Cognitive Combination Principles
Synthesis Templates for In-depth Thinking.
Recall that in Section 3.1, we generate $k$ single-hop seed questions and proactively inform $\pi$ that these questions will be used for future cognitive combinations. Here, we explicitly specify some cognitive combination relationships into the synthesis requirement $P_{task}$ for multi-hop QAs, and let $\pi$ to self-challenging by figuring out how to combine the generated $\{q^{1}_{i},...q^{k}_{i}\}_{i\in N}$ into more complex questions $\{Q_{i}\}_{i\in N}$ conditioned on aforementioned $P_{cog}$ and $P_{context}$, such that the results is logically right, cognitively considered and match the most suitable relationships among the given ones.
Building on the single-hop seed questions generated in Section 3.1, we advance to multi-hop question synthesis by explicitly defining cognitive combination relationships within the task specification $P_{\text{task}}$. During this phase, the model $\pi$ is prompted to synthesize more complex questions $\{Q_{i}\}_{i\in N}$ by logically combining the pre-generated seed questions $\{q^{j}_{i}\}_{i=1,...,N;j=1,...,k}$ under constraints from $P_{\text{cog}}$ (cognitive framework) and $P_{\text{context}}$ (factual reference). This process aims to make $\{Q_{i}\}_{i\in N}$ logically valid, cognitively grounded, and optimally match the most relevant relationships from our predefined taxonomy.
Textual Cognitive Combination Types.
For $P_{\text{context}}$ containing only textual content, we define three cognitive combination forms after initial seed generation:
  • Comparison: Requires contrasting two or more concepts or scenarios.
  • Bridging: Connects disparate concepts through intermediate reasoning steps.
  • Temporal: Involves analyzing events across chronological contexts.
Multimodal Cognitive Combination Types.
For $P_{\text{context}}$ containing both text and images, we introduce four multimodal combination types:
  • Visual-Textual Alignment: Ensures consistency between visual data (e.g., diagrams, charts) and textual descriptions.
  • Spatial Reasoning: Addresses spatial relationships or geometric arrangements depicted in visual content.
  • Causal Inference: Requires deducing cause-effect relationships from combined textual and visual inputs.
  • Contextual Synthesis: Demands integration of information across modalities (e.g., text, images, graphs) to form unified conclusions.
3.2.2 Cognitive Manifestation
Structured Cognitive Processes & Reasoning Units.
Beyond specifying cognitive combination relationships in the task requirement $P_{\text{task}}$, we explicitly require $\pi$ to articulate its intermediate thought process $T_{i}$ leading to the final question $Q_{i}$. This support $T_{i}$ links the seed questions $\{q^{j}_{i}\}_{i\in N}$ into a logically coherent reasoning pathway, explicitly connecting steps to ensure interpretability and contextual consistency.
To balance cognitive context relevance and inference efficiency, $\pi$ is prompted to jointly generate $Q_{i}$, $A_{i}$, and $T_{i}$ in one inference step. However, raw responses may include fragmented or overly lengthy content, so we further prompt $\pi$ to segment outputs into distinct components: formal question-answer pairs $(Q_{i},A_{i})$ and their supporting rationale $T_{i}$. This results in modular cognitive units of varying granularity, enabling more flexible refinement of $\pi$’s reasoning abilities in subsequent training (more details are in later Section 3.3).
Figure 2: An end-to-end example of MindGYM. For multimodal data, we generate five seed questions first, and make the model self-challenging itself via synthesizing multi-hop questions and multi-hop answers while preserving its internal thinking process.
Cognitive Information Processing Dynamics.
Note that the synthesis framework explicitly involves $\pi$ processing both novel and existing knowledge during inference. By reintroducing context elements from $P_{\text{context}}$, accessing seed questions $\{q^{j}_{i}\}_{i\in N}$, and addressing advanced combination requirements in $P_{\text{task}}$, the model is incentivized to: (1). Explore diverse reasoning patterns embedded in $P_{\text{cog}}$, (2). Re-examine context details with renewed emphasis, and (3). Construct multi-hop answers using enriched relational constraints. This process yields more focused, nuanced questions $Q_{i}$ while eliminating reasoning shortcuts, forcing $\pi$ to engage in deeper inferential analysis.
Figure 2 illustrates an example in which the original $P_{\text{context}}$ (including image and textual inputs) is first used to synthesize seed questions (e.g., “What ingredients did Kenji use?” or “What vitamins are in broccoli?”), as detailed in Section 3.1). These are then combined via the multi-hop synthesis protocol (outlined in Section 3.2) into more challenging questions that require cross-contextual reasoning (e.g., “How do Kenji’s lunch ingredients compare with Tracy’s in terms of vitamin content?” paired with bridging and temporal relationships).
Critically, whereas the original QA focused on limited details (e.g., Kenji’s tomatoes and Tracy’s broccoli), our synthesized questions expand reasoning depth by incorporating broader contextual elements: sandwich compositions, citrus fruits, root vegetables, etc. The resulting multi-hop QAs demonstrate richer interconnections, broader scope, and higher cognitive demands on the solver.
3.3 Thinking-Induced Curriculum Fine-Tuning
Training Objectives.
Given the self-synthesized dataset containing triples $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$, we propose a “MindGym”-style training flow to progressively integrate cognitive guidance into the model $\pi$ through mixed masking strategies. This approach incrementally challenges $\pi$ by varying the exposure to question-answer pairs ($Q_{i}$, $A_{i}$), their supporting rationales ($T_{i}$), and masked input-output configurations.
For recap, $Q_{i}$ represents the complex multi-hop question, $A_{i}$ denotes the reference answer with structured reasoning, and $T_{i}$ records the detailed step-by-step cognitive trajectory. We construct four training datasets with distinct masking objectives:
$\mathcal{D}^{1}_{\text{train}}=\{(Q_{i},T_{i}\rightarrow A_{i})\}$ employs the masked input structure $\langle Q_{i},T_{i},[\mathbf{\text{MASK}}]\rangle$. The model predicts $A_{i}$ using $Q_{i}$ and explicit thinking guidance $T_{i}$, and the loss is defined as:
$\mathcal{L}_{1}=-\mathbb{E}_{(Q,T,A)\sim\mathcal{D}^{1}_{\text{train}}}\log\pi_{\theta}(A|Q,T).$ (1)
$\mathcal{D}^{2}_{\text{train}}=\{(Q_{i},A_{i}\rightarrow T_{i})\}$ uses the structure $\langle Q_{i},[\mathbf{\text{MASK}}],A_{i}\rangle$. The objective is to recover the hidden thinking path $T_{i}$, with loss:
$\mathcal{L}_{2}=-\mathbb{E}_{(Q,A,T)\sim\mathcal{D}^{2}_{\text{train}}}\log\pi_{\theta}(T|Q,A).$ (2)
$\mathcal{D}^{3}_{\text{train}}=\{Q_{i}\rightarrow(A_{i},T_{i})\}$ follows $\langle Q_{i},[\mathbf{\text{MASK}}]\rangle$, requiring simultaneous generation of $A_{i}$ and $T_{i}$:
$\mathcal{L}_{3}=-\mathbb{E}_{(Q,A,T)\sim\mathcal{D}^{3}_{\text{train}}}\log\pi_{\theta}(A,T|Q).$ (3)
$\mathcal{D}^{4}_{\text{train}}=\{Q_{i}\rightarrow A_{i}\}$ uses the same masking strategy as $\mathcal{D}_{train}^{3}$, with input $\langle Q_{i},[\mathbf{\text{MASK}}]\rangle$. This mimics standard supervised fine-tuning practice:
$\mathcal{L}_{4}=-\mathbb{E}_{(Q,A)\sim\mathcal{D}^{4}_{\text{train}}}\log\pi_{\theta}(A|Q).$ (4)
The curriculum progresses through these datasets sequentially: $\mathcal{D}^{1}\rightarrow\mathcal{D}^{2}\rightarrow\mathcal{D}^{3}\rightarrow\mathcal{D}^{4}$. This mirrors a guided learning pathway where $\pi$ first reproduces answers with external cognitive scaffolding ($\mathcal{D}^{1}$), then deduces thinking paths without full guidance ($\mathcal{D}^{2}$), gradually internalizes end-to-end reasoning ($\mathcal{D}^{3}$), and finally masters standalone Q-to-A mappings ($\mathcal{D}^{4}$). Empirical validation in Sec. 4.3 shows this strategy outperforms non-curriculum mixed training.
Discussion on Efficiency and Cost.
The MindGYM framework enhances $\pi$’s reasoning capabilities through $N\times 4$ synthesis inferences across stages ($P_{\text{context}}$, seed questions, $Q_{i}$, and $\mathcal{D}_{\text{train}}$). Despite this incremental setup, inference efficiency is maintained through amortized reasoning costs during training. A single inference suffices to achieve high model performance, as shown empirically in Section 4.3. Unlike test-time scaling approaches [38, 18], we move cognitive enhancement costs to the training phase. Compared to data-intensive methods (e.g., reinforcement learning) requiring extensive trial-and-error [19], our method maximizes cognitive depth leveraging structured synthetic data without resource-heavy exploration phases.
Datasets Multimodal Eval Text-Only Eval Avg
MMStar MathVista MathVision MM-Avg GSM8K MATH GPQA Text-Avg
Baslines Raw 64.00 69.30 24.67 52.66 83.62 67.60 31.83 61.02 56.84
LIMO (817) [43] 63.93 70.20 25.9 53.37 84.08 67.80 30.58 60.82 57.10
Open-O1 (77k) [35] 58.93 61.40 17.43 45.92 77.94 59.00 28.19 55.04 50.48
Open-O1 (400) [35] 63.87 69.70 23.66 53.08 84.31 69.00 29.7 61.00 57.04
MMEvol-SciQA (106) [24] 63.93 69.50 23.36 52.26 83.40 65.60 31.83 60.28 56.27
MMEvol-DVQA (4k) [24] 62.93 67.40 24.01 51.45 83.85 65.80 33.08 60.91 56.18
Text MindGYM (400) 64.33 70.30 28.62 54.42 84.08 68.40 31.33 61.27 57.84
Image MindGYM-OKVQA (400) 63.60 69.20 28.29 53.70 84.00 68.20 31.08 61.09 57.40
MindGYM-SciQA (400) 64.20 69.80 25.33 53.11 83.93 69.00 31.45 61.46 57.29
Table 1: Performance of Qwen2.5-VL-7B on various evaluation benchmarks. The number following the dataset name indicates its size in terms of the number of samples used for training.
4 Experiments
To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative text and multimodal evaluation sets, comparing it with several competitive data synthesis methods and datasets. Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix A.
4.1 Experimental Settings
Datasets.
We conduct multi-modality data synthesis across textual and image contexts. For textual synthesis, data generation occurs exclusively through self-contained knowledge extraction from the base model, eliminating external data dependencies. Multimodal synthesis utilizes established scientific and reasoning corpora as visual primitives, with visual grounding from ScienceQA [22] and OK-VQA [25].
Benchmarks.
To systematically assess MindGYM on VLMs across reasoning modalities, we incorporate:
  • Multimodal evaluation sets include MMStar [9], MathVista-Mini [23], and MathVision-Mini [37], designed to comprehensively assess the reasoning abilities in the multimodal scenario.
  • Text-based evaluation sets comprise GSM8K [10], MATH [15], and GPQA [31], focused on evaluating reasoning and mathematical capabilities.
The MM-Avg and Text-Avg metrics serve as unified criteria for assessing multimodal and text-based reasoning capabilities respectively, whereas Avg quantifies comprehensive performance spanning both unimodal and multimodal reasoning domains.
Models & Implementation.
The data synthesis of MindGYM leverages VLMs’ inherent long-context understanding and reasoning capabilities. We implement Qwen2.5-VL-7B-Instruct [4] as our base VLM since alternative models (e.g., InternVL2.5-7B) exhibit inferior generation quality under identical configurations (detailed in Appendix D). To maintain linguistic consistency with Qwen2.5-VL-7B-Instruct and consider the information density of reasoning data, we utilize Chinese for data synthesis. All experiments follow consistent protocols: during text-only training, only LLM-layers update parameters, while image-text joint training additionally activates Aligner-layers, keeping ViT-layers permanently frozen.
Baselines.
As MindGYM generates self-contained adversarial training data without external sources, direct comparison with conventional data synthesis methods proves infeasible. We therefore select SOTA reasoning-oriented dataset works from two categories: (1) Multimodal reasoning using MMEvol’s [24] core subsets ScienceQA and DVQA, and (2) Text-based reasoning with LIMO [43], 817 curated logical chains, and Open-O1 [35], an SFT dataset for CoT activation as fundamental text reasoning baselines.
4.2 Main Results
To validate the efficacy of our methodology, we systematically evaluate the model across multiple benchmarks as documented in Table 1, with key insights summarized:
Universal Superiority.
MindGYM demonstrates substantial performance advantages through comprehensive average metric evaluation in MindGYM-Text with a 54.42 MM-Avg-score and 57.84 Avg-score, and notably achieving 28.62 on MathVision with a 16% improvement over the Raw model. The proposed method significantly outperforms all baselines where even the strongest competitor LIMO trails by 1.05 points in MM-Avg. An interesting phenomenon emerges among baselines that Open-O1 (77K) suffers severe degradation with 6.74-point MM-Avg decline and 29.3% MathVision collapse despite employing a large number of training samples, exposing the vulnerability of conventional scaling approaches.
Cross-modal Transfer.
MindGYM enables effective cross-modal reasoning capability sharing through bidirectional knowledge transfer. Remarkably, the text-specialized MindGYM-Text establishes new SOTA performance on MM-Avg, with the score of 54.42 that surpasses all baselines. Conversely, the multimodal-enhanced MindGYM-SciQA demonstrates unexpected supremacy in textual reasoning tasks, achieving an optimal Text-Avg score of 61.46. This bidirectional enhancement confirms the transferability of reasoning capabilities between textual and multimodal representations.
Generalizability & Robustness.
Consistent improvements emerge across diverse data sources. Our approach attains sub-optimal average performance when synthesizing data from both SciQA, showing an improvement of 0.45 over raw data, and OKVQA, with an improvement of 0.56 over raw data, demonstrating adaptability to various domain-specific reasoning tasks. This indicates that our method effectively enhances generalizability and robustness across diverse reasoning challenges.
Data Efficiency.
MindGYM actives model potential using only 400 synthesized samples, achieving a 28.62 MathVision score that surpasses baselines trained with 10$\times$-100$\times$ more data. This contrasts sharply with Open-O1’s performance collapse despite massive data scaling, indicating that MindGYM significantly enhances the model’s reasoning capabilities while also ensuring that the model’s performance is not detrimentally affected.
These findings effectively validate that our approach addresses the issues of modality inconsistency, generalization, and computational efficiency mentioned in Section 2.
4.3 Ablation Studies
To validate the efficacy of our approach, we conduct systematic module-wise utility analysis through targeted ablation studies.
GPT-based Thinking Quality Scoring.
Datasets Depth Breadth Avg
MindGYM/ Raw 10.2% 19.4% 14.8% $\uparrow$
MindGYM/ LIMO 1.63% 23.8% 12.7% $\uparrow$
MindGYM/ Open-O1 (400) 1.88% 23.7% 12.8% $\uparrow$
MindGYM/ MMEvol-DVQA 8.41% 37.0% 22.7% $\uparrow$
Overall 5.53% $\uparrow$ 26.0% $\uparrow$ 15.8% $\uparrow$
Table 2: The results of win rates in terms of relative improvements over baselines, using GPT-4 as a scorer.
We establish a two-dimensional evaluation framework leveraging GPT-4 as an expert scorer. The protocol assesses: (1) Reasoning Depth via derivation complexity, and (2) Reasoning Breadth by solution diversity. Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix B.5. Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark. The experimental outcomes documented in Table 2 demonstrate MindGYM’ consistent superiority across both evaluation axes, particularly achieving 15.8% average improvements in reasoning depth over the raw model and other competitive baselines.
Datasets MM-Avg Text-Avg Avg
Raw 52.66 61.02 56.84
MindGYM-Text 54.42 61.27 57.84
MindGYM-Text w/o SC 52.72 60.00 56.36 $\downarrow$
MindGYM-Text Syn-EN 52.77 60.30 56.53 $\downarrow$
MindGYM-Text w/o CF 52.20 60.83 56.51 $\downarrow$
MindGYM-Text with RB 52.20 60.83 56.51 $\downarrow$
MindGYM-OKVQA 53.70 61.09 57.40
MindGYM-OKVQA w/o SC 52.85 61.30 57.08 $\downarrow$
MindGYM-OKVQA Syn-EN 53.14 60.59 56.87 $\downarrow$
MindGYM-OKVQA with RB 53.33 60.60 56.97 $\downarrow$
MindGYM-SciQA 53.11 61.46 57.29
MindGYM-SciQA w/o SC 53.20 60.63 56.91 $\downarrow$
MindGYM-SciQA Syn-EN 52.25 60.88 56.56 $\downarrow$
MindGYM-SciQA with RB 53.46 60.27 56.87 $\downarrow$
Table 3: Ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of Curriculum Fine-tuning (w/o CF) steps and Relation Balanced (with RB), complete results detailed in Appendix C.
Ablation of Stream of Consciousness.
We conduct the ablation study by removing the stream of consciousness module from our prompt while maintaining the same curriculum fine-tuning method for training and evaluation. The experimental results, as shown in Table 3, indicate that the removal of the stream of consciousness module leads to a decrease across all Avg scores, with reductions ranging from 0.32 to 1.48 points. This decline highlights the significance of the stream of consciousness module in our approach, underscoring its essential role in enhancing the depth of reasoning and overall performance of the model.
Curriculum Fine-tuning Order.
To verify the effectiveness of our Curriculum Fine-tuning module, we mix and shuffle the first three datasets ($\mathcal{D}^{1}_{\text{train}}$, $\mathcal{D}^{2}_{\text{train}}$, $\mathcal{D}^{3}_{\text{train}}$) and train the model on text data, subsequently using $\mathcal{D}^{4}_{\text{train}}$ to constrain the answers’ generation pattern. This experimental setup allows us to systematically examine whether disrupting the curriculum order impacts model performance. The experimental results shown in Table 3 under the w/o CF condition reveal a performance degradation of 1.33 points, highlighting the critical impact of the Curriculum Fine-tuning order on model performance.
Linguistic Difference.
Based on the selected Qwen series [4] model, we utilize Chinese as the language for data synthesis. Additionally, we explore the effects of using English for synthesis, as shown in Table 3 under the Syn-EN condition. The results indicate varying degrees of performance degradation when using English, with a decrease of up to 1.31 points in the MindGYM-Text setting. Beyond the inherent characteristics of Qwen, we hypothesize that the higher information density of Chinese makes it more beneficial for the model’s reasoning learning.
Relationship Distribution Balanced.
To verify the effectiveness of allowing the model to autonomously select relationship categories, we balance the number of entries generated for each category, totaling 400 entries. The experimental results shown in Table 3 under the with RB condition indicate that enforcing category balance leads to a decline up to 1.33 in overall model performance. This finding suggests that the autonomous selection of relationships by the model reflects its cognitive understanding.
5 Conclusion & Future Works
This work introduces MindGYM, a framework that enhances vision-language reasoning via synthetic self-challenging questions. By self-challenging multi-step reasoning data and employing structured curriculum training, MindGYM achieves strong performance gains with small data size and reduced computational costs, outperforming prior methods in reasoning depth and breadth.
Future directions include extending cognitive relationships to dynamic visual scenarios [46]; exploring the application of MindGYM in specific domains like medicine [27] and finance; developing adaptive scoring operators [6] for question complexity; and integrating agents [13] for data correctness verification in generalized RL environments.
Appendix A Implementation of Experiments
A.1 Utilized Seed Visual Datasets
For the multimodal domain, we leverage widely-adopted scientific and reasoning datasets as visual source material for question synthesis, specifically including ScienceQA and OK-VQA.
  • ScienceQA [22], the first large-scale multimodal dataset that annotates lectures and explanations for the answers.
  • OK-VQA [25], a dataset for visual question answering that requires methods which can draw upon outside knowledge to answer questions.
A.2 Benchmarks
To evaluate the models’ reasoning capabilities and performance in both multimodal and text comprehension, we adopt the methodology used by major advanced open-source VLM Qwen2.5-VL-7B [4], selecting the following widely utilized reasoning evaluation datasets. All evaluations were conducted on the ms-swift [45] platform.
  • Multimodal Benchmarks:
    • MMStar [9], an elite vision-indispensable multi-modal benchmark comprising 1,500 challenge samples meticulously selected by humans.
    • MathVista (Mini) [23], a benchmark designed to combine challenges from diverse mathematical and visual tasks, consists of 6,141 examples, derived from 28 existing multimodal datasets.
    • MathVision (Mini) [37], a meticulously curated collection of 3,040 high-quality mathematical problems with visual contexts sourced from real math competitions.
  • Text-based Benchmarks:
    • GSM8K [10], a dataset of high quality linguistically diverse grade school math word problems created by human problem writers, which takes between 2 and 8 steps to solve.
    • MATH [15], a dataset of challenging competition mathematics problems, each problem in MATH has a full step-by-step solution.
    • GPQA [31], stands for Graduate-Level Google-Proof Q&A Benchmark, a challenging dataset designed to evaluate the capabilities of LLMs and scalable oversight mechanisms.
Given the extensive evaluation tasks, employing the complete evaluation sets would incur substantial time expenditure. To expedite the evaluation process while ensuring fairness and accuracy, we adopt the $eval\_limit$ parameter from ms-swift, configuring GPQA to 300 samples, while retaining the original sample sizes for the remaining benchmark datasets. All experiments were conducted using this consistent setup to ensure the fairness of the experiments.
A.3 Training and Evaluation Details
Platform
We implement our approaches using PyTorch [29] v2.5.1, coupled with PEFT v0.14.0 and the Transformers library [41] v4.49.0. Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with LLMs loaded as 16-bit floating-point numbers. The specific data-model development processes are completed in Data-Juicer Sandbox [8], via integration with the ms-swift [45] repository for training and evaluation, and the VLMEvalKit [11] repository for evaluation.
Baselines
Our method requires no external background knowledge for text generation and enhances reasoning capabilities through the model’s inherent visual understanding. While this approach differs fundamentally from typical data synthesis methodologies, we select representative reasoning-enhancing datasets for comparative analysis to validate the effectiveness of our approach in SFT for VLMs.
  • Multimodal Reasoning:
    • MMEvol[24]:A method addressing data quality limitations by generating complex and diverse image-text instruction datasets to improve VLMs. We consider two subsets of MMEvol: (1) For direct comparison, we utilize its ScienceQA synthetic data subset. (2) For comprehensive reasoning comparison, we adopt DVQA which primarily focusing on mathematical and logical reasoning tasks.
  • Text-Based Reasoning:
    • LIMO [43]: Challenges conventional assumptions in mathematical reasoning by demonstrating that models achieve superior performance with smaller quantities of high-quality training data. We reformat its ”question+solution+answer” structure into standard instruction pairs for training consistency.
    • Open-O1 [35]: An open-source initiative to replicate the reasoning capabilities of proprietary models through curated SFT data for CoT activation. The complete 77k-sample dataset trains both LLaMA and Qwen architectures. To evaluate data efficiency, we additionally create a 400-sample randomly subsampled version for performance validation on same scale.
Training Details
In our experimental setup, we employ Low-Rank Adaptation (LoRA) [16] adapters for the fine-tuning process, utilizing a LoRA-rank of 8 and a LoRA-alpha of 32. The learning rate was consistently maintained at $1\times 10^{-5}$ on Qwen2.5-VL-7B-Instruct. We utilize a batch size of 4 and set the maximum sequence length to 4096 tokens to accommodate the model’s reasoning capacity. To optimize the training process, a warmup ratio of 0.05 was applied and a validation ratio of 0.03 was used. The training was conducted over a single epoch, balancing computational efficiency with the need for effective model adaptation.
In the VLM training configuration, for pure textual information, we freeze both the Vision Transformer (ViT) and the Aligner while keeping the LLM layers active, to ensure that the image processing capabilities of the VLM remain unaffected. For image-text data, we adopt the post-tuning strategy of the Qwen2.5-VL series [4], where the ViT remains frozen while both the Aligner and LLM are activated to maximize multimodal information acquisition.
Evaluation Details
Following the evaluation protocols established in ms-swift [45] and VLMEvalKit [11], we maintained default configurations with pt as the inference backend. For multimodal tasks, we implemented VLMEvalKit as the evaluation backend, while employing the Native framework for text-only evaluation scenarios.
Appendix B Prompts
B.1 Stream of Consciousness
The protocol introduce in Section 3.1 is based on Thinking Claude, which is an innovative prompt engineering framework designed to enhance the reasoning capabilities of Claude AI models through simulated human-like cognitive processes. This paradigm-shifting approach employs a multi-stage cognitive architecture featuring chain-of-thought prompting, self-reflective mechanisms, and structured problem decomposition. We incorporate it into MindGYM’s data synthesis process to enhance the inference performance of the model [^2] https://github.com/richards199999/Thinking-Claude.
B.2 Prompts for Background Generation
On the text side, due to the low quality of sub-question generation without foundational documents, we create background documents for sub-question generation through iterative polling based on eight cognitive reasoning themes. On the image side, since text-to-image generation models are still underdeveloped and produce low-quality results, we directly utilize existing reasoning and trial-based QA datasets.
[…]
The document should be rich enough to support multi-hop reasoning in Chinese.
B.3 Prompts for Seed Single-Hop Question Synthesis
These two prompts are used to synthesize sub-problems, Prompt 2 is used to synthesize text sub-problems, and Prompt 3 is used to synthesize images sub-problems.
[…]
Based on the background document provided, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to a single category:
- Bridging: requiring connecting facts or pieces of information from the document.
- Comparison: involving comparing two or more elements described in the document.
- Temporal: requiring reasoning about the order or timing of events.
Clearly state the relationship category that links all sub-questions in Chinese:
[…]
Based on the provided image, original question, and original answer, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to one of the following categories:
- Visual-Textual Alignment: Requiring alignment between visual (e.g., images, charts) and textual information.
- Spatial Reasoning: Involving spatial relationships or geometric layouts.
- Causal Inference: Requiring reasoning about cause-and-effect relationships.
- Contextual Synthesis: Requiring synthesis of information across multiple modalities (e.g., text, images, charts).
Clearly state the relationship category that links all sub-questions in Chinese:
B.4 Prompts for Challenging Multi-Hop QA Synthesis
These two prompts are used to synthesize multi-hop problems, Prompt 4 is used to synthesize multi-hop problems of text, and Prompt 5 is used to synthesize multi-hop problems of images.
[…]
Synthesize a multi-hop question and its answer based on the above sub-questions and background document in Chinese. Please start Qwen thinking and return the thinking process:
[…]
The question should require reasoning across the sub-questions and synthesizing information from the image and original question.
Then, provide a detailed answer to the multi-hop question, ensuring it is consistent with the image and sub-questions.
Synthesize a multi-hop question and its answer based on the above sub-questions, original image, and original question in Chinese. Please start Qwen thinking and return the thinking process:
B.5 Prompts for GPT-based Scorer
This is a prompt scored with GPT4 to evaluate the scores of the two models in terms of depth and breadth.
[…]
As a mathematical problem solver, please strictly compare and analyze the predictions of the two models:
[Problem description]
{question}
[Original model prediction]
{pred_raw}
[Fine-tuned model prediction]
{pred_finetuned}
Evaluation requirements:
Give the scores of raw and finetuned respectively
1. Reasoning depth score (0-3 points):
- 3 points: multi-step derivation with verification
- 2 points: complete derivation steps
- 1 point: simple calculation steps
- 0 points: no derivation process
2. Reasoning breadth score (0-3):
- 3 points: Explores multiple valid methods/angles, justifies optimal choice
- 2 points: Mentions alternative approaches briefly but focuses on one
- 1 points: Suggests another method without analysis
- 0 points: Single approach with no alternatives considered
3. Comparative analysis: use bullet points to list the main improvements/degradations
Please return a JSON that strictly follows this format: ”accuracy_raw”: 0-2, ”accuracy_finetuned”: 0-2, ”depth_raw”: 0-3, ”depth_finetuned”: 0-3, ”comparison”: [”Point 1”, ”Point 2”]
Appendix C Complete Results including Ablation Study
We present the full experimental results in Table 4, comprehensively validating the effectiveness of our method compared to other baselines, as well as the robustness of our approach itself (ablation studies introduced in Section 4.3).
Datasets Multimodal Eval Text-Only Eval Avg
MMStar MathVista MathVision MM-Avg GSM8K MATH GPQA Text-Avg
Baslines Raw 64.00 69.30 24.67 52.66 83.62 67.60 31.83 61.02 56.84
LIMO (817) [43] 63.93 70.20 25.90 53.37 84.08 67.80 30.58 60.82 57.10
Open-O1 (77k) [35] 58.93 61.40 17.43 45.92 77.94 59.00 28.19 55.04 50.48
Open-O1 (400) [35] 63.87 69.70 23.66 53.08 84.31 69.00 29.70 61.00 57.04
MMEvol-SciQA (106) [24] 63.93 69.50 23.36 52.26 83.40 65.60 31.83 60.28 56.27
MMEvol-DVQA (4k) [24] 62.93 67.40 24.01 51.45 83.85 65.80 33.08 60.91 56.18
Text MindGYM 64.33 70.30 28.62 54.42 84.08 68.40 31.33 61.27 57.84
MindGYM w/o SC 63.60 69.90 24.67 52.72 83.83 66.60 29.57 60.00 56.36
MindGYM Syn-EN 63.93 69.70 24.67 52.77 84.00 67.20 29.70 60.30 56.53
MindGYM w/o CF 64.00 69.90 22.69 52.20 84.15 67.00 31.33 60.83 56.51
MindGYM w RB 64.00 70.10 22.70 52.20 84.15 67.00 31.33 60.83 56.51
Image MindGYM-OKVQA 63.60 69.20 28.29 53.70 84.00 68.20 31.08 61.09 57.40
MindGYM-OKVQA w/o SC 63.93 69.30 25.33 52.85 84.15 68.80 30.95 61.30 57.08
MindGYM-OKVQA Syn-EN 63.53 69.90 25.99 53.14 84.23 67.60 29.95 60.59 56.87
MindGYM-OKVQA w RB 63.80 71.20 25.00 53.33 84.00 66.60 31.20 60.60 56.97
MindGYM-SciQA 64.20 69.80 25.33 53.11 83.93 69.00 31.45 61.46 57.29
MindGYM-SciQA w/o SC 64.07 69.20 26.32 53.20 83.93 67.00 30.95 60.63 56.91
MindGYM-SciQA Syn-EN 63.73 69.00 24.01 52.25 83.55 68.00 31.08 60.88 56.56
MindGYM-SciQA w RB 64.20 70.20 25.99 53.46 83.85 65.00 31.96 60.27 56.87
Table 4: Complete ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of Curriculum Fine-tuning (w/o CF) steps and Relation Balanced (with RB).
Appendix D Alternative VLMs Data Synthesis through MindGYM
As demonstrated in Listing 1, our analysis using InternVL2.5-8B [5] reveals its generated outputs tend to exhibit chaotic patterns. We hypothesize this stems from dual factors: (i) constrained input token limits restricting contextual comprehension, and (ii) inherent deficiencies in cognitive reasoning capabilities. These limitations necessitated its exclusion from our comparative VLM experiments.
Listing 1: Data Sample of MindGYM on InternVL2.5-8B.
1{
2 "background_document": "Background document for multi-hop reasoning.",
3 "reasoning_category": "Mathematical Reasoning",
4 "sub_questions": [
5 "What is the optimal pricing strategy considering cost inflation?",
6 "How will fixed and variable costs vary with seasonal demand?",
7 "What are the impacts of geopolitical changes on supply chain?",
8 "What role do regulatory changes play in pricing?",
9 "How will budget allocation differ among different product lines?"
10 ],
11 "relationship_category": "Comparison",
12 "multihop_question": "What steps need to be taken to optimize pricing strategies taking into account rising costs, shifting regulations, seasonal fluctuations, and differing product revenues?",
13 "multihop_answer": "To maximize profit under increased costs, companies should reassess fixed and variable expenses, implement strategic cost savings wherever feasible, monitor regulatory developments closely, review existing budgets, allocate resources accordingly, update pricing schedules, communicate revisions internally consistently, measure effectivity quarterly.",
[…]
15}